OpenStudy (anonymous):

Can somebody please explain to me how 15 divided by 30% equals 50. I dont understand the steps.

OpenStudy (anonymous):

I need to understand the steps.

OpenStudy (accessdenied):

It might help to understand what 30% is as a real number. The percentage itself doesn't give us much meaning, so we should understand that it is 30 parts per 100.

OpenStudy (anonymous):

?

OpenStudy (accessdenied):

Can you rewrite 30% as a decimal or fraction? This is what I am getting at. :p

OpenStudy (anonymous):

Wouldnt you have to change 30% to .3

OpenStudy (accessdenied):

Yes. Now you should be able to divide these two numbers. You could use a calculator, or multiply numerator and denominator by 10 to get rid of decimals. \( \dfrac{15}{.3} \times \dfrac{10}{10} = \dfrac{150}{3} \)

OpenStudy (anonymous):

I am sorry, i am really bad at math. I dont comprehend where you got 10

OpenStudy (lovelyharmonics):

that has to be one of the scariest things ive ever seen... ۞_۞ sorry i couldn't resist telling you that

OpenStudy (lovelyharmonics):

* your picture i mean bye! c:

OpenStudy (anonymous):

I am not here to be judged help or get out

OpenStudy (accessdenied):

If we want to move the decimal point in 0.3, we have to multiply by 10's because the first decimal place represents "tenths". And we can multiply both numerator and denominator by 10 simultaneously because we are effectively multiplying by 1 (any number divided by itself is 1).

OpenStudy (anonymous):

k

OpenStudy (anonymous):

i still dont see what your point is

OpenStudy (accessdenied):

15/.3 would be pretty difficult to divide on its own. If you can do so, then you should outright find it is equal to 50. 15 = 3*5 = 0.3 * 50, dividing 0.3. The purpose of multiplying both numerator and denominator by 10 is because 150 / 3 can be more easily divided (working with whole numbers is generally nicer).

OpenStudy (anonymous):